Given that HCF (306, 657) = 9, find LCM (306, 657).

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

As we know that,

HCF × LCM = Product of the two given numbers

Therefore,

9 × LCM = 306 × 657

LCM = (306 × 657)/9 = 22338

Hence, LCM(306,657) = 22338

Thanks